OpenStudy (anonymous):

-27 greater than or equal to 3x-9 How do you solve this? Taking a refresher course and need some help to jog my memory

OpenStudy (tfht):

add nine to both sides to get -18>or=to 3x then divide both sides by three to get -6>or= to x

OpenStudy (anonymous):

add 9 to what? the 3x and the -9 or the -27?

OpenStudy (anonymous):

He said...."to both sides" .. The whole idea in algebra is that you always do the same thing to both sides to keep the equality.

OpenStudy (texaschic101):

-27 >= 3x - 9 --- add 9 to both sides -27 + 9 >= 3x - 9 + 9 -- combine like terms - 18 >= 3x -- divide both sides by 3 -18/3 >= 3x/3 -6 > x do you understand ?

OpenStudy (anonymous):

Thankyou @texaschic101 you made it more clear to understand

OpenStudy (texaschic101):

glad to help :)